my homework shows a table with 3 fractions, 1 over 8, 1 over 3, and 1 over 7
now the headline says "reciprocals: copy and complete the table" and I have no
idea about what is going on, can you help me?

Found 2 solutions by checkley77, drk:
Answer by checkley77(12569)

RECIPROCAL OF A FRACTION IS THE INVERSE OF THE FRACTION.
RECIPROCAL OF 1/8=8/1=8
RECIPROCAL OF 1/3=3/1=3
RECIPROCAL 0F 1/7=7/1=7

Answer by drk(1908)

When you say copy and complete the table, the table might have headings of fraction, decimal, percent.
FRACTION . . . . . . . . . . . . .DECIMAL . . . . . . . . . . . . .PERCENT
(1/8) . . . . . . . . . . . . . . . . .(0.125) . . . . . . . . . . . . . . .(12.5%)
(1/3). . . . . . . . . . . . . . . . .(0.333) . . . . . . . . . . . . . . .(33.33%)
(1/7). . . . . . . . . . . . . . . . .(0.142857) . . . . . . . . . . . . (14 and 2/7%)
maybe you mean this??